# FM4Com: Foundation Model for Scene-Adaptive Communication Strategy Optimization

### v1

@@ Abstract
The emergence of sixth-generation (6G) networks heralds an intelligent communication ecosystem driven by AI-native air interfaces. However, current physical-layer designs—typically following modular and isolated optimization paradigms—fail to achieve global end-to-end optimality due to neglected inter-module dependencies. Although large language models (LLMs) have recently been applied to communication tasks such as beam prediction and resource allocation, existing studies remain limited to single-task or single-modality scenarios and lack the ability to jointly reason over communication states and user intents for personalized strategy adaptation. To address these limitations, this paper proposes a novel multimodal communication decision-making model based on reinforcement learning. The proposed model semantically aligns channel state information (CSI) and textual user instructions, enabling comprehensive understanding of both physical-layer conditions and communication intents. It then generates physically realizable, user-customized link construction strategies that dynamically adapt to changing environments and preference tendencies. A two-stage reinforcement learning framework is employed: the first stage expands the experience pool via heuristic exploration and behavior cloning to obtain a near-optimal initialization, while the second stage fine-tunes the model through multi-objective reinforcement learning considering bit error rate, throughput, and complexity. Experimental results demonstrate that the proposed model significantly outperforms conventional planning-based algorithms under challenging channel conditions, achieving robust, efficient, and personalized 6G link construction.
@@ I Introduction
To address the aforementioned challenges, we propose a foundation model designed for interactive communication strategy customization with a Chain-of-Thought-enhanced Reinforcement Learning (CoT-RL) framework named IntraCom-FM, which enables the construction of end-to-end communication links that adapt to specific user preferences under varying channel conditions. The main contributions of this paper can be summarized as follows:
  • We propose a foundation model for scenario-adaptive communication strategy optimization, termed FM4Com. The model possesses the capability to comprehensively perceive both CSI and user intents. By semantically integrating physical-layer conditions with natural language instructions, FM4Com can generate customized and optimal link configuration strategies that align with the current channel dynamics and user preferences, thereby achieving personalized and adaptive optimization of communication systems.
  • We propose a chain-of-thought-enhanced two-stage reinforcement learning framework. In the first stage, a heuristic exploration algorithm is employed to expand the reinforcement learning experience pool, followed by behavior cloning via supervised learning to obtain a near-optimal initialization. In the second stage, the model is fine-tuned through reinforcement learning to further improve decision quality. Throughout the training process, the model also generates user preference representations as intermediate reasoning steps, forming an internal chain of thought that guides subsequent decision-making.
  • Experimental results demonstrate that the proposed FM4Com model can adaptively construct robust and efficient link strategies tailored to diverse user requirements under various challenging channel conditions, significantly outperforming existing planning-based algorithms.
@@ II-A Full-link Communication Simulation System
We consider a 3GPP-compliant physical-layer communication link[16], as illustrated in Fig. 1. The system consists of several functional modules, including channel coding, spreading, modulation, power control, resource allocation, channel estimation, equalization, and HARQ. Each module adopts a specific strategy denoted as $c_{i}$, and the combination of all selected strategies forms a complete physical-layer transmission chain $\bm{a}=\{a_{1},a_{2},\ldots,a_{n}\}$.
@@
Based on the relationship between transmitted and received data, the overall performance of the physical-layer system can be comprehensively evaluated in terms of BER, system rate, and complexity.
@@ III Proposed Method
In this section, we introduce the structure of our proposed FM4Com and the CoT-RL training method.
@@ III-A Multimodal Communication Decision Large Language Model
As illustrated in Fig. 1, the FM4Com consists of a text encoder, a CSI encoder, a cross-modal attention module for textual semantic compression, a pretrained LLM, and a generator responsible for final strategy generation. The generator comprises $N_{a}$ actor networks, each designed to produce the optimal configuration strategy for a specific physical-layer module.
Figure 1: Our proposed FM4Com.
@@
$\bm{{\mathit{Y}}}_{\text{csi}}=\text{Encoder}_{\text{csi}}(\bm{H})),$ (4)
@@
$\bm{T}_{\text{text}}=\text{Tokenizer}(\bm{T}),$ (5)
where $\bm{T}_{\text{text}}\in\mathbb{R}^{L_{1}\times d}$ represents the word embeddeing of the original text.
@@
$\tilde{\bm{Z}}_{\text{text}}=\text{Self-Attention}(\tilde{\bm{Z}}_{\text{text}}),$ (7)
@@
where $\bm{{\mathit{Y}}}_{\text{LLM}}\in\mathbb{R}^{2L\times d}$ denotes the hidden representation from the final transformer layer, and $L_{c}$ is the length of the compressed text sequence. Within the reasoning process, the LLM generates an intermediate preference vector $\bm{{\mathit{p}}}$ that reflects the user’s optimization tendency, such as throughput-oriented or complexity-sensitive objectives.
Finally, the preference vector $\bm{{\mathit{p}}}$ is concatenated with the multimodal LLM output and passed through multiple actor networks to produce the final decision strategies for different physical-layer modules. The output of the $i$-th actor network can be expressed as
$a_{i}=g_{\text{actor},i}\!\left(\text{Concat}[\bm{{\mathit{Y}}}_{\text{LLM}};\bm{{\mathit{p}}}]\right),\quad i=1,2,\ldots,N_{a},$ (11)
where $N_{a}$ denotes the number of decision modules such as channel coding, modulation, power control, and resource allocation. The overall physical-layer configuration is thus represented as
@@
III-B CoT-RL Training Method
Figure 2: RL method.
The proposed model is trained using a two-stage CoT-RL procedure, which integrates behavior cloning for initialization and reinforcement learning for fine-tuning. The entire training process aims to optimize the multimodal decision model such that it can reason over both textual and CSI features to generate adaptive and high-performance communication strategies.
In the first stage, we perform behavior cloning to pretrain the model using a set of high-quality samples collected through a greedy search algorithm. Specifically, during early exploration, the multimodal model interacts with the environment to generate decision samples $(\bm{{\mathit{s}}},\bm{{\mathit{a}}},r)$, where $\bm{{\mathit{s}}}$ denotes the state composed of textual and CSI embeddings, $\bm{{\mathit{a}}}$ represents the selected strategy $\bm{{\mathit{a}}}$, and $r$ is the obtained reward. The greedy algorithm selects actions that yield the highest immediate reward, and the resulting tuples are stored in a replay buffer $\mathcal{D}$. The model is then trained to imitate these expert-like decisions through supervised learning:
$\mathcal{L}_{\text{BC}}=-\mathbb{E}_{(\bm{{\mathit{s}}},\bm{{\mathit{a}}})\sim\mathcal{D}}\big[\log\pi_{\theta}(\bm{{\mathit{a}}}|\bm{{\mathit{s}}})\big],$ (13)
where $\pi_{\theta}$ denotes the policy parameterized by the LLM-based actor and $\theta$ represents the trainable parameters. This behavior cloning stage enables the model to quickly learn a near-optimal policy by mimicking the expert trajectories in the buffer, effectively providing a warm start for subsequent reinforcement learning.
After behavior cloning, the model undergoes reinforcement learning fine-tuning to further optimize decision performance under the true environment dynamics. Given the state $\bm{{\mathit{s}}}$, the model samples an action $\bm{{\mathit{a}}}\sim\pi_{\theta}(\bm{{\mathit{a}}}|\bm{{\mathit{s}}})$, receives a scalar reward $r=f_{\text{reward}}(\text{BER},\text{Rate},\text{Complexity})$, and updates its parameters to maximize the expected return:
@@ III-B CoT-RL Training Method
To stabilize training, the behavior cloning loss is combined with the reinforcement learning objective to form the final joint loss:
$\mathcal{L}_{\text{total}}=-J(\theta)+\lambda\mathcal{L}_{\text{BC}},$ (16)
where $\lambda$ is a balancing coefficient controlling the influence of imitation learning on policy optimization.
@@
$r=w_{\text{ber}}\cdot R_{\text{ber}}+w_{\text{rate}}\cdot R_{\text{rate}}+w_{\text{comp}}\cdot R_{\text{comp}},$ (17)
where $w_{\text{ber}},w_{\text{rate}},w_{\text{comp}}\in[0,1]$ are weighting coefficients satisfying $w_{\text{ber}}+w_{\text{rate}}+w_{\text{comp}}=1$, and the normalized terms $R_{\text{ber}}$, $R_{\text{rate}}$, and $R_{\text{comp}}$ represent the contributions of BER, rate, and complexity, respectively.
To reflect user-specific communication preferences, these weights are adaptively determined based on the preference vector $\bm{{\mathit{p}}}$ inferred by the LLM:
$w_{\text{ber}},w_{\text{rate}},w_{\text{comp}}]=f_{\text{pref}}(\bm{{\mathit{p}}}),$ (18)
where users emphasizing reliability are assigned larger $w_{\text{ber}}$, while those focusing on high throughput or low complexity receive higher $w_{\text{rate}}$ or $w_{\text{comp}}$, respectively. Moreover, during the entire training process, we incorporate a cross-entropy loss $\mathcal{L}_{\text{pref}}$ on user preference prediction into the overall objective to ensure that the model accurately understands the user intent.
@@ IV Numerical Results
This section presents our simulation results. First, we describe the simulation setup in detail. Then, we compare FM4Com with several existing policy selection methods to evaluate the effectiveness of the proposed approach. In addition, we present a human-computer interactive question-answering example to demonstrate the intelligence and flexibility of FM4Com.
@@ IV-A Experimental Setup
We employ the QuaDRiGa channel generator to create a time-varying CSI dataset compliant with the 3GPP standard, covering multiple scenarios such as urban, rural, and highway environments. Corresponding textual descriptions are generated based on the characteristics of each scenario. To represent user preferences, we randomly assign three types of preference settings across all scenarios: low bit error rate, high throughput, and conventional systems. The GPT-2 model is adopted as the pretrained LLM backbone for validation.
@@ IV-B Model Decision Performance
We first analyze the FM4Com’s ability to generate personalized communication strategies under different user preferences. For comparison, we employ three baseline methods for strategy selection: random strategy selection, greedy search, and beam search, with the beam width of the beam search method set to 3. As illustrated in Fig.3, the proposed model and baseline methods are evaluated across various channel environments in terms of BER, system complexity, and system rate.
@@
(c) Complexity
Figure 3: Performance comparison of different strategy selection methods for low BER strategy under variou SNRs.
For the low-BER preference, the FM4Com effectively captures the user’s intent and outputs robust communication strategies. As the SNR increases, the BER of all methods decreases; however, the reduction achieved by random selection is minimal, while the greedy and beam search algorithms demonstrate moderate improvements. In contrast, FM4Com significantly outperforms all baselines, achieving near-zero BER across a wide SNR range. This advantage arises because greedy and beam search methods only optimize local module-level performance. In contrast, FM4Com leverages reinforcement learning with end-to-end feedback, enabling globally optimal decisions across the entire communication chain. It is also observed that the remarkably low BER achieved by the proposed model comes at the cost of reduced throughput, which aligns with the user’s low-BER preference emphasizing reliability over efficiency.
@@
(c) Complexity
Figure 4: Performance comparison of different strategy selection methods for high rate strategy under various SNRs.
For the high-throughput preference, the FM4Com adapts its strategy accordingly, achieving higher throughput than in the low-BER case while maintaining a lower BER than random selection. Although greedy and beam search methods can yield higher throughput, their solutions are overly aggressive, leading to extremely high BER and rendering the constructed communication links nearly incapable of transmitting valid information. In contrast, our model achieves a better trade-off between throughput and reliability, balancing user preference with practical system stability.
Under the low-complexity preference, the FM4Com tends to adopt more conservative strategies, exhibiting similar behavior to the low-BER preference scenario. This is because low-BER configurations are inherently stable across different environments, and the model naturally gravitates toward such reliable solutions when computational simplicity is prioritized.
@@
(c) Complexity
Figure 5: Performance comparison of different strategy selection methods for general strategy under various SNRs.
@@ IV-C Model Complexity
Random 0.004s
FM4Com 0.46 s
Beam Search 36.36s
Greedy 13.55s
To assess the practical feasibility of deploying the proposed framework in real-world communication systems, we further analyze the algorithmic complexity of different strategy planning methods in a single planning task. As shown in Table I, heuristic-based algorithms such as the greedy and beam search methods exhibit significantly higher inference latency compared with the proposed FM4Com. This is primarily due to their sequential search and evaluation nature, where each potential action combination must be explicitly explored or locally optimized, resulting in exponential or at least polynomial growth in computational cost as the action space increases. In contrast, our FM4Com leverages end-to-end learning to implicitly capture the complex dependencies between modules, allowing it to generate globally optimized strategies through a single forward pass of the neural network. This substantial reduction in inference time demonstrates the scalability and practicality of the proposed model.
@@ IV-D Model Interaction
Figure 6: FM4Com interaction examples.
In this section, we present a concrete human–machine interaction example to demonstrate the functional capabilities and intelligent characteristics of the proposed FM4Com. As illustrated in Fig.6, under a given communication condition, the user provides both the known environmental information and the specific transmission intent as multimodal inputs to the model.Upon receiving these inputs, FM4Com accurately interprets the user’s intent, comprehends the underlying communication context, and infers the user’s preference orientation. For example, according to the query shown in Fig.6, the model can infer that the user emphasizes reliability in the communication system. Based on this understanding, the model autonomously generates an optimized and interpretable communication strategy configuration, which aligns with both the current channel state and the user’s semantic preference. This example highlights the model’s strong semantic reasoning and adaptive decision-making capabilities, demonstrating its potential to serve as an intelligent agent for human–AI collaborative communication system design.

### v2

# Agentic Link Construction for Environment and Intent Aware 6G Communication

@@ Abstract
The emergence of sixth-generation (6G) networks heralds an intelligent communication ecosystem driven by the rapid proliferation of intelligent services and increasingly complex communication scenarios. However, current physical-layer designs—typically following modular and isolated optimization paradigms—fail to achieve global end-to-end optimality due to neglected inter-module dependencies. Although large language models (LLMs) have recently been applied to communication tasks such as beam prediction and resource allocation, existing studies remain limited to single-task or single-modality scenarios and lack the ability to jointly reason over communication states and user intents for personalized strategy adaptation. To address these limitations, this paper proposes a novel multimodal communication decision-making model for link construction leveraging reinforcement learning on pretrained LLMs. The proposed model semantically aligns channel state information (CSI) and textual user instructions, enabling comprehensive understanding of both physical-layer conditions and communication intents. It then generates physically realizable, user-customized link construction that dynamically adapts to changing environments and preference tendencies. A two-stage reinforcement learning framework is employed: the first stage expands the experience pool via heuristic exploration and behavior cloning to obtain a near-optimal initialization, while the second stage fine-tunes the model through multi-objective reinforcement learning considering BER, throughput, and power consumption. Experimental results demonstrate that the proposed model significantly outperforms conventional planning-based algorithms under challenging channel conditions, achieving robust, efficient, and personalized end-to-end communication strategies. Our dataset and code are available at https://github.com/LZhaoyang/AGENS.
@@ I Introduction
To address the aforementioned challenges, we propose a foundation model designed for interactive communication strategy customization with Proximal Policy Optimization (PPO) reinforcement learning framework named AGENS, which enables the construction of end-to-end communication links that adapt to specific user preferences under varying channel conditions. The main contributions of this paper can be summarized as follows:
  • We propose a foundation model for scenario-adaptive communication strategy optimization, termed AGENS. The model possesses the capability to comprehensively perceive both CSI and user intents. By semantically integrating physical-layer conditions with natural language instructions, AGENS can generate customized and optimal link configuration strategies that align with the current channel dynamics and user preferences, thereby achieving personalized and adaptive optimization of communication systems.
  • We propose a two-stage reinforcement learning framework. In the first stage, a heuristic exploration algorithm is employed to expand the reinforcement learning experience pool, followed by behavior cloning via Supervised Fine-Tuning (SFT) to obtain a near-optimal initialization. In the second stage, the model is fine-tuned through reinforcement learning to further improve decision quality.
  • Experimental results demonstrate that the proposed AGENS model can adaptively construct robust and efficient link strategies tailored to diverse user requirements under various challenging channel conditions, significantly outperforming existing planning-based algorithms.
@@ II-A Full-link Communication Simulation System
We consider a 3GPP-compliant physical-layer communication link[16], as illustrated in Fig. 1. The system consists of several functional modules, including channel coding, coding rate selection, modulation scheme, precoding, power control, channel estimation, and equalization. Each module adopts a specific strategy denoted as $c_{i}$, and the combination of all selected strategies forms a complete physical-layer transmission chain $\bm{a}=\{a_{1},a_{2},\ldots,a_{n}\}$.
@@
Based on the relationship between transmitted and received data, the overall performance of the physical-layer system can be comprehensively evaluated in terms of BER, system rate, and power consumption.
@@ III Proposed Method
In this section, we introduce the structure of our proposed AGENS and the training method.
@@ III-A Multimodal Communication Decision Large Language Model
As illustrated in Fig. 1, the AGENS consists of a text encoder, a CSI encoder, a cross-modal attention module for textual semantic compression, a pretrained LLM, and a generator responsible for final strategy generation. The generator comprises $n$ actor networks, each designed to produce the optimal configuration strategy for a specific physical-layer module.
Figure 1: Our proposed AGENS.
@@
$\bm{{\mathit{Y}}}_{\text{csi}}=\text{Encoder}_{\text{csi}}(\bm{H}),$ (4)
@@
$\bm{T}_{\text{text}}=\text{Embed}(\text{Tokenizer}(\bm{T})),$ (5)
where $\bm{T}_{\text{text}}\in\mathbb{R}^{L_{1}\times d}$ represents the word embedding of the original text.
@@
$\tilde{\bm{Z}}_{\text{text}}=\text{Self-Attention}({\bm{Z}}_{\text{text}}),$ (7)
@@
where $\bm{{\mathit{Y}}}_{\text{LLM}}\in\mathbb{R}^{2L\times d}$ denotes the hidden representation from the final transformer layer.
Finally, the multimodal feature $\bm{\mathit{Y}}_{\mathrm{LLM}}$ is fed into multiple executor networks to generate the final decision strategies for different physical-layer modules. The output of the $i$-th executor network can be expressed as:
$a_{i}=g_{\text{actor},i}\!\left(\bm{{\mathit{Y}}}_{\text{LLM}}\right),\quad i=1,2,\ldots,n,$ (11)
where $n$ denotes the number of decision modules such as channel coding, modulation, power control, and resource allocation. The overall physical-layer configuration is thus represented as:
@@
III-B Training Method
Figure 2: Illustration of the proposed RL method.
As illustrated in Fig. 2, the proposed model is trained using a two-stage procedure that combines SFT and reinforcement learning fine-tuning. The objective of this training strategy is to optimize the multimodal decision model so that it can effectively reason over both textual and CSI features to generate adaptive and high-performance communication strategies.
In the first stage, SFT is conducted using a set of high-quality decision samples generated by a greedy search algorithm. During early exploration, the multimodal model interacts with the environment to produce decision tuples $(\bm{\mathit{s}},\bm{\mathit{a}},r)$, where $\bm{\mathit{s}}$ denotes the state consisting of textual and CSI embeddings, $\bm{\mathit{a}}$ represents the selected communication strategy, and $r$ is the corresponding reward. The greedy search algorithm selects actions that maximize the immediate reward, and the resulting tuples are stored in a replay buffer $\mathcal{D}$. The model is then pretrained in a supervised manner to imitate these high-quality decisions by minimizing the discrepancy between the predicted actions and the stored actions in $\mathcal{D}$ :
$\mathcal{L}_{\text{SFT}}=-\mathbb{E}_{(\bm{{\mathit{s}}},\bm{{\mathit{a}}})\sim\mathcal{D}}\big[\log\pi_{\theta}(\bm{{\mathit{a}}}|\bm{{\mathit{s}}})\big],$ (13)
where $\pi_{\theta}$ denotes the policy parameterized by the LLM-based actor and $\theta$ represents the trainable parameters. This SFT stage enables the model to quickly learn a near-optimal policy by mimicking the expert trajectories in the buffer, effectively providing a warm start for subsequent reinforcement learning.
After SFT, the model undergoes reinforcement learning fine-tuning to further optimize decision performance under the true environment dynamics. Given the state $\bm{{\mathit{s}}}$, the model samples an action $\bm{{\mathit{a}}}\sim\pi_{\theta}(\bm{{\mathit{a}}}|\bm{{\mathit{s}}})$, receives a scalar reward $r=f_{\text{reward}}(\text{BER},\text{Rate},\text{Power})$, and updates its parameters to maximize the expected return:
@@ III-B Training Method
$r=w_{\text{ber}}\cdot R_{\text{ber}}+w_{\text{rate}}\cdot R_{\text{rate}}+w_{\text{power}}\cdot R_{\text{power}},$ (16)
where $w_{\text{ber}},w_{\text{rate}},w_{\text{power}}\in[0,1]$ are weighting coefficients satisfying $w_{\text{ber}}+w_{\text{rate}}+w_{\text{power}}=1$, and the normalized terms $R_{\text{ber}}$, $R_{\text{rate}}$, and $R_{\text{power}}$ represent the contributions of BER, rate, and power consumption, respectively.
@@ IV Numerical Results
This section presents our simulation results. First, we describe the simulation setup in detail. Then, we compare AGENS with several existing policy selection methods. In addition, we present a human-computer interactive question-answering example to demonstrate the intelligence and flexibility of AGENS.
@@ IV-A Experimental Setup
We generate CSI datasets using Sionna-RT with diverse parameter configurations. For each CSI sample, a corresponding natural-language user-intent label is assigned to build diverse task scenarios. In the design of user intents, we consider three representative 6G service preferences: high throughput, high reliability, and energy-efficient operation. To evaluate the performance of the constructed communication link, we adopt BER, rate (defined as the ratio of the number of successfully transmitted bits to the number of original bits), and power compensation (i.e., the amount of transmit-side compensation power, measured in dB) as the evaluation metrics. For model implementation, GPT-2 Medium (355M parameters) is employed as the backbone architecture for building the AGENS. It should be noted that the pretrained LLM is mainly employed for its strong language understanding and user-intent inference capabilities. The proposed framework can be easily replaced with other pretrained language models.
@@ IV-B Model Decision Performance
We first evaluate the capability of AGENS to generate personalized communication strategies under varying user preferences. For comparison, we consider four baseline methods: random strategy selection, greedy search, beam search (with a beam width of 2), and a Transformer-based variant of AGENS, where the LLM component is replaced by a standard Transformer. The Transformer-based baseline is introduced to highlight the effectiveness of incorporating LLMs within our framework. For the random strategy baseline, different random seeds are employed for different user preferences to ensure sufficient diversity in the generated strategies.
@@
(c) Power Consumption
Figure 3: Performance comparison of different strategy selection methods for high throughput strategy under various SNRs.
As shown in Fig. 3, when high throughput is prioritized, all methods trade BER for higher transmission rates. Beam search achieves the best performance among heuristic baselines due to its larger search space, yet remains limited by local optimization. In contrast, AGENS attains higher rates while maintaining BER within an acceptable range, achieving a better balance between throughput and reliability. When reliability is emphasized, AGENS generates more robust strategies that better align with user intent. As illustrated in Fig. 4, although both heuristic methods and AGENS produce low-BER strategies, AGENS shows clear advantages in low-SNR regimes. This is partly because it adopts lower transmission rates and higher transmit power, consistent with practical system design principles. Under the energy efficiency objective, beam search and AGENS achieve comparable performance. However, AGENS still attains slightly lower BER while maintaining low power consumption, indicating its superior global planning capability over heuristic methods.
@@
(c) Power Consumption
Figure 4: Performance comparison of different strategy selection methods for high reliability strategy under various SNRs.
It can also be observed that the Transformer-based method performs worse than traditional heuristic approaches when either high throughput or high reliability is prioritized. Although it achieves relatively low power consumption under the energy efficiency objective, it often produces overly aggressive strategies. As shown in Fig. 5, the Transformer-based method tends to generate rigid decisions, selecting nearly identical transmission rates across all scenarios. Such strategies result in excessively high BER, which may render the communication system unreliable in practice. These observations highlight the necessity of incorporating LLMs. With stronger reasoning and contextual understanding capabilities, LLMs enable more effective and adaptive communication strategy planning.
Finally, some performance curves exhibit noticeable fluctuations. This is partly due to the inherent randomness of the random strategy, and partly because heuristic methods rely on local optimization of individual modules, which may lead to unstable overall performance in multi-module communication systems. In contrast, AGENS produces smoother curves across different metrics, indicating better global coordination and higher stability.
@@
(c) Power Consumption
Figure 5: Performance comparison of different strategy selection methods for energy-efficient strategy under various SNRs.
@@ IV-C Model Complexity
Random 0.005s
AGENS 0.167 s
Beam 446.472s
Greedy 307.987s
To evaluate the practical feasibility of deploying the proposed framework in real-world communication systems, we analyze the algorithmic complexity of different strategy planning methods for a single planning task. As shown in Table I, heuristic-based methods such as greedy search and beam search exhibit significantly higher inference latency than AGENS due to their sequential search and evaluation processes, where candidate actions must be explicitly explored or optimized. In contrast, AGENS leverages end-to-end learning to capture the dependencies among modules and generate globally optimized strategies through a single forward pass. This greatly reduces inference time and demonstrates the scalability and practicality of the proposed model.
@@ IV-D Model Interaction
Figure 6: An input-output example of AGENS.
In this section, we present a human–machine interaction example to illustrate the capabilities of the proposed AGENS. As shown in Fig. 6, under a given communication condition, the user provides environmental information and transmission intent as multimodal inputs to the model. Based on these inputs, AGENS interprets the user intent, understands the communication context, and infers the user preference. For example, from the query in Fig. 6, the model identifies that high throughput is prioritized in the communication system. Accordingly, it generates an optimized and interpretable communication strategy.